# LLMs as In-Context Meta-Learners for Model and Hyperparameter Selection

### v1

@@ 1 Introduction
The performance of machine learning (ML) models hinges on the selection of appropriate algorithms and their hyperparameters. This joint optimization task is commonly referred to as the Combined Algorithm Selection and Hyperparameter optimization (CASH) problem (Thornton et al., 2013; Bergstra and Bengio, 2012; Snoek et al., 2012). Traditionally, practitioners have relied on manual tuning, grid search, or Bayesian optimization techniques (Mockus et al., 1978; Shahriari et al., 2016) to navigate this complex search space. However, these approaches are computationally expensive and demand substantial domain expertise. This creates barriers to entry and limits the scalability of ML applications across diverse domains.
Large language models (LLMs) have recently shown strong capabilities in reasoning, knowledge synthesis, and problem-solving across domains (Wei et al., 2022). As they scale, they exhibit emergent behaviors that enable adaptation to new tasks by reusing prior experience in context (Brown et al., 2020; Dong et al., 2024). These behaviors have been interpreted as a form of in-context meta-learning, with transformers proposed as general-purpose meta-learners (Kirsch et al., 2024) and LLMs studied explicitly in this role (Coda-Forno et al., 2023). Much of this prior work has focused on demonstrating the phenomenon itself, often in synthetic or language-oriented tasks. By contrast, model and hyperparameter selection provides a practical and consequential setting in machine learning where generalization across tasks directly impacts performance and efficiency. If LLMs can transfer knowledge in this context, they may offer a new paradigm for addressing the CASH problem and extend our understanding of their adaptability beyond controlled demonstrations.
@@
This research introduces two prompting strategies for leveraging LLMs in model and hyperparameter selection. The Zero-Shot strategy relies solely on high-level task metadata, requiring no prior examples. The Meta-Informed strategy augments this by incorporating pairs of task metadata and well-performing model configurations from previous tasks, enabling more informed recommendations (Figure 1). Unlike prior work (Zheng et al., 2023; Zhang et al., 2024), our approach operates without iterative validation feedback. It also enables cross-task generalization in the meta-informed case. Importantly, we prompt the LLM to propose complete configurations consisting of both model families and associated hyperparameters, which can then be directly evaluated or integrated into downstream pipelines.
We evaluate both prompting strategies on tabular regression and classification tasks. Results show that LLMs, when properly prompted, can make surprisingly effective recommendations even in zero-shot settings where conventional methods often require extensive experimentation. The meta-informed strategy further improves performance by leveraging prior knowledge, often approaching or matching the quality of expert-guided selections. Taken together, these findings highlight the potential of LLMs as meta-learners in automated machine learning: they can reason about datasets, models, and hyperparameters with minimal tuning, offering a scalable and accessible alternative to traditional search-based or expert-driven workflows. This also complements concurrent applications of LLMs to other stages of the AutoML pipeline such as feature engineering with CAAFE (Hollmann et al., 2023).
@@ 2 Related Work
Early work on hyperparameter optimization (HPO) relied on simple search strategies such as grid search and random search (Bergstra and Bengio, 2012). More sophisticated model-based methods, such as Bayesian optimization (BO), iteratively fit surrogate models to past evaluations and propose promising configurations (Bergstra et al., 2011; Snoek et al., 2012). Subsequent advances introduced multi-fidelity and bandit-based approaches, including Successive Halving (Jamieson and Talwalkar, 2016) and Hyperband (Li et al., 2017), which exploit early stopping to allocate resources efficiently. Later extensions sought to transfer knowledge across related tasks or account for computational budgets, for example through multi-task Bayesian optimization and compute-aware methods (Swersky et al., 2013; Golovin et al., 2017). These methods significantly improved efficiency but still treat each optimization task largely in isolation.
@@
To overcome this limitation, meta-learning approaches aim to accelerate HPO by leveraging prior experience across tasks. Transfer Neural Processes (TNP) (Wei et al., 2021), for example, incorporate meta-knowledge such as surrogate models and historical trial data to improve sample efficiency. Meta-Bayesian optimization methods extend this idea by learning priors over surrogate models from related tasks, enabling faster convergence on new optimization problems (Feurer et al., 2015; Perrone et al., 2018). Other approaches, such as ALFA (Baik et al., 2020), adapt hyperparameters dynamically during training using a meta-learner, while SHSR (Borboudakis et al., 2023) prunes unpromising regions of the search space using past AutoML runs. PriorBand (Mallik et al., 2023) further accelerates HPO by combining expert beliefs with low-fidelity proxy tasks to guide search in deep learning pipelines. These methods illustrate the value of meta-knowledge, but they still assume a fixed model class.
@@
In practice, algorithms and hyperparameters must be optimized jointly, formalized as the CASH problem (Thornton et al., 2013). A common approach is to treat model choice as a categorical hyperparameter, as in Auto-WEKA (Thornton et al., 2013) and Auto-sklearn (Feurer et al., 2015), but the resulting search space is large and expensive to explore. Bandit-based formulations address this by casting algorithm selection as arms with HPO inside each arm, e.g., MaxUCB (Balef et al., 2025), Rising Bandits (Li et al., 2020), and ER-UCB (Hu et al., 2021). These improve scalability but still depend on extensive search. In contrast, our method tackles CASH directly by generating model and hyperparameter configurations without relying on hierarchical search or bandit-style exploration.
@@
LLMs have recently been applied to hyperparameter optimization, for example through iterative refinement with feedback or by combining with Bayesian optimization (Zhang et al., 2024; Mahammadli and Ertekin, 2025; Liu et al., 2025). While promising, these approaches treat HPO in isolation and require multiple interaction rounds. By contrast, we address the broader CASH problem, producing complete model–hyperparameter configurations in a single inference. AutoML-GPT (Zhang et al., 2023) explores full pipeline automation, including preprocessing, but depends on explicit task similarity matching. Our method is simpler and more practical: we use prior tasks only as in-context examples, letting the LLM adapt implicitly, and we evaluate directly on real-world tabular datasets under standard CASH protocols.
@@ 4 Motivation: Synthetic Ridge Regression Experiment
We evaluate the Qwen 2.5 family (7B, 14B, 32B, 72B) (Qwen et al., 2025), across decoding temperatures $\{0.0,0.2,0.4,0.6,0.8\}$. Prompt templates are provided in Appendix A.3. To ensure valid outputs, generations are limited to 5 tokens with invalid predictions resampled.
@@ 5.1 Methodology
In practice, the Meta-Informed strategy assumes access to previous tasks along with high-performing configurations. For this study, we obtained such configurations by running extensive hyperparameter search with HEBO (Cowen-Rivers et al., 2022) on a set of tabular regression and classification tasks. To maximize performance, ensembles (or blends) were built from the resulting models. We refer to the models with the highest contributions to these ensembles as Context Blends, and use them as the source of support examples passed to the prompt.
@@
The LLM is instructed to output a JSON object describing an ensemble of 10 models. For each supported family (CatBoost (Prokhorenkova et al., 2018), LightGBM, XGBoost (Chen and Guestrin, 2016), and scikit-learn MLP (Pedregosa et al., 2011)), we provide the model name, a list of valid hyperparameters, and a discrete grid of admissible values. This grid is included directly in the prompt, ensuring that the model generates configurations from a well-defined search space rather than free-form values. An excerpt of the schema is shown below (see Appendix F for full hyperparamater grids):
@@
We use the DeepSeek-R1 reasoning model (DeepSeek-AI et al., 2025), which naturally produces explanations of its choices. The LLM configuration is described in Appendix D. Invalid generations are rare, but we apply lightweight post-processing when they occur. If the LLM outputs a numeric value that falls outside the predefined hyperparameter grid, we project it to the nearest valid grid point. For non-numeric fields (e.g., categorical options) that cannot be matched, we discard the configuration and resample a fresh output. Likewise, if the JSON structure itself is malformed, the entire configuration is rejected and regenerated. Each run uses a different set of support examples, ensuring robustness to contextual variation.
@@
Each LLM call outputs 10 configurations, which we treat as candidate base models. We train these with cross-validation bagging and then combine their predictions using feedforward greedy blending (Caruana et al., 2004). This procedure is applied consistently to LLM-based and baseline methods, providing a fair comparison and reflecting common ML ensembling practice.
@@ 5.3 Baselines
We compare LLM-based recommendations against four baselines representing different strategies for the CASH problem (full details in Appendix G): Context-Random (uniformly samples model–hyperparameter configurations from the same reference pool as the one passed to the LLM), Random-Hyperopt (at each step, uniformly samples a model family and then applies a hyperparameter optimizer within that family), LGBM-Hyperopt (optimizer restricted to LightGBM, capturing the strength of a single tuned family), and MaxUCB-Hyperopt (treats each family as a bandit arm, selecting the one with the highest upper-confidence bound before a single optimization step (Balef et al., 2025)). Context Blends consist of ensembles obtained from an extensive hyperparameter search. They provide upper-bound baselines: they achieve high performance through extensive search, and thus set the performance we seek to approach under a much more limited budget. All -Hyperopt baselines use HEBO (Cowen-Rivers et al., 2022), chosen for its strong and consistent performance across diverse tasks (Kegl, 2023) [^1] HEBO begins with random search, using $1+\text{(dimension of the hyperparameter space)}$ evaluations, before switching to Bayesian optimization..
@@ A.1 Closed-Form Test Error
    Deterministic equivalents / resolvent fixed point. Using standard random-matrix techniques (resolvent identities and deterministic equivalents for sample covariance resolvents) [Couillet and Debbah, 2011, Chapter 6], one shows that the random matrix inverse that appears in the ridge formula concentrates around the deterministic matrix $\bar{\mathbf{Q}}(\lambda)$ defined in equation 4 and that the scalar traces $(1/n)\operatorname{tr}(\mathbf{C}_{k}\bar{\mathbf{Q}})$ converge to the solution $\delta_{k}$ of equation 5. This gives the first-order deterministic equivalents used to compute $m_{k}$.
@@
    Gaussian fluctuation / CLT. After centering by the deterministic mean, the score is a linear or quadratic form of Gaussian vectors; a multivariate CLT (together with second-order deterministic equivalents captured by $\mathbf{K}_{k}$ and the $\mathbf{d}^{(j)}$ corrections) yields asymptotic Gaussianity Tiomoko et al. [2020] with variance given by the deterministic formula $v_{k}$.
@@ Appendix F Base Model Details
We use four base models in our experiments: XGBoost [Chen and Guestrin, 2016], CatBoost [Prokhorenkova et al., 2018], LGBM, and SKMLP [Pedregosa et al., 2011]. The corresponding hyperparameter grids for each model are provided in Figure 6.
@@ G.4 MaxUCB-Hyperopt
For the MaxUCB-Hyperopt baseline, we implement the bandit-based CASH formulation proposed by Balef et al. [2025]. In this setting, each candidate model family is treated as an arm in a multi-armed bandit, and hyperparameter optimization is carried out within the selected arm using HEBO. The Max-UCB algorithm balances exploration of new model families with exploitation of those that have already demonstrated promising performance.
@@ Appendix H Robustness to Prompt Shuffling
Large language models can sometimes exhibit position or recency biases [Wang et al., 2023, 2025], raising the question of whether the Meta-Informed strategy is sensitive to the way information is ordered inside the prompt. To test this, we generate two independent shuffled versions of the Meta-Informed prompt for each dataset–seed pair. In each shuffle, we randomly permute (i) the order of support datasets, (ii) the order of model families listed in the schema, and (iii) the order of hyperparameters within each family. The underlying content is unchanged, only the presentation order differs. The experimental setup is otherwise identical on the 22 Kaggle datasets the same contexts, ensembling pipeline, and $p_{\text{rank}}$ as the evaluation metric.

### v2

@@ 1 Introduction
The performance of machine learning (ML) models hinges on the selection of appropriate algorithms and their hyperparameters. This joint optimization task is commonly referred to as the Combined Algorithm Selection and Hyperparameter optimization (CASH) problem (thornton2013auto; bergstra2012; Snoek2012). Traditionally, practitioners have relied on manual tuning, grid search, or Bayesian optimization techniques (Mockus1978; Shahriari2016TakingTH) to navigate this complex search space. However, these approaches are computationally expensive and demand substantial domain expertise. This creates barriers to entry and limits the scalability of ML applications across diverse domains.
Large language models (LLMs) have recently shown strong capabilities in reasoning, knowledge synthesis, and problem-solving across domains (Wei2022). As they scale, they exhibit emergent behaviors that enable adaptation to new tasks by reusing prior experience in context (Brown2020; dong-etal-2024-survey). These behaviors have been interpreted as a form of in-context meta-learning, with transformers proposed as general-purpose meta-learners (kirsch2024generalpurposeincontextlearningmetalearning) and LLMs studied explicitly in this role (coda-forno2023metaincontext). Much of this prior work has focused on demonstrating the phenomenon itself, often in synthetic or language-oriented tasks. By contrast, model and hyperparameter selection provides a practical and consequential setting in machine learning where generalization across tasks directly impacts performance and efficiency. If LLMs can transfer knowledge in this context, they may offer a new paradigm for addressing the CASH problem and extend our understanding of their adaptability beyond controlled demonstrations.
@@
This research introduces two prompting strategies for leveraging LLMs in model and hyperparameter selection. The Zero-Shot strategy relies solely on high-level task metadata, requiring no prior examples. The Meta-Informed strategy augments this by incorporating pairs of task metadata and well-performing model configurations from previous tasks, enabling more informed recommendations (Figure 1). Unlike prior work (Zheng2023; zhang2024usinglargelanguagemodels), our approach operates without iterative validation feedback. It also enables cross-task generalization in the meta-informed case. Importantly, we prompt the LLM to propose complete configurations consisting of both model families and associated hyperparameters, which can then be directly evaluated or integrated into downstream pipelines.
We evaluate both prompting strategies on tabular regression and classification tasks. Results show that LLMs, when properly prompted, can make surprisingly effective recommendations even in zero-shot settings where conventional methods often require extensive experimentation. The meta-informed strategy further improves performance by leveraging prior knowledge, often approaching or matching the quality of expert-guided selections. Taken together, these findings highlight the potential of LLMs as meta-learners in automated machine learning: they can reason about datasets, models, and hyperparameters with minimal tuning, offering a scalable and accessible alternative to traditional search-based or expert-driven workflows. This also complements concurrent applications of LLMs to other stages of the AutoML pipeline such as feature engineering with CAAFE (hollmann2023large).
@@ 2 Related Work
Early work on hyperparameter optimization (HPO) relied on simple search strategies such as grid search and random search (bergstra2012). More sophisticated model-based methods, such as Bayesian optimization (BO), iteratively fit surrogate models to past evaluations and propose promising configurations (Bergstra2011; Snoek2012). Subsequent advances introduced multi-fidelity and bandit-based approaches, including Successive Halving (Jamieson2016) and Hyperband (Li2017), which exploit early stopping to allocate resources efficiently. Later extensions sought to transfer knowledge across related tasks or account for computational budgets, for example through multi-task Bayesian optimization and compute-aware methods (Swersky2013; Golovin2017). These methods significantly improved efficiency but still treat each optimization task largely in isolation.
@@
To overcome this limitation, meta-learning approaches aim to accelerate HPO by leveraging prior experience across tasks. Transfer Neural Processes (TNP) (Wei2021), for example, incorporate meta-knowledge such as surrogate models and historical trial data to improve sample efficiency. Meta-Bayesian optimization methods extend this idea by learning priors over surrogate models from related tasks, enabling faster convergence on new optimization problems (Feurer2015; Perrone2018). Other approaches, such as ALFA (Baik2020), adapt hyperparameters dynamically during training using a meta-learner, while SHSR (borboudakis2023meta) prunes unpromising regions of the search space using past AutoML runs. PriorBand (mallik2023priorband) further accelerates HPO by combining expert beliefs with low-fidelity proxy tasks to guide search in deep learning pipelines. These methods illustrate the value of meta-knowledge, but they still assume a fixed model class.
@@
In practice, algorithms and hyperparameters must be optimized jointly, formalized as the CASH problem (thornton2013auto). A common approach is to treat model choice as a categorical hyperparameter, as in Auto-WEKA (thornton2013auto) and Auto-sklearn (Feurer2015), but the resulting search space is large and expensive to explore. Bandit-based formulations address this by casting algorithm selection as arms with HPO inside each arm, e.g., MaxUCB (balef2025cashbanditsmaxkarmed), Rising Bandits (Li2020), and ER-UCB (Hu2021). These improve scalability but still depend on extensive search. In contrast, our method tackles CASH directly by generating model and hyperparameter configurations without relying on hierarchical search or bandit-style exploration.
@@
LLMs have recently been applied to hyperparameter optimization, for example through iterative refinement with feedback or by combining with Bayesian optimization (zhang2024usinglargelanguagemodels; mahammadli2025sequentiallargelanguagemodelbased; liu2025agenthpo). While promising, these approaches treat HPO in isolation and require multiple interaction rounds. By contrast, we address the broader CASH problem, producing complete model–hyperparameter configurations in a single inference. AutoML-GPT (zhang2023automlgptautomaticmachinelearning) explores full pipeline automation, including preprocessing, but depends on explicit task similarity matching. Our method is simpler and more practical: we use prior tasks only as in-context examples, letting the LLM adapt implicitly, and we evaluate directly on real-world tabular datasets under standard CASH protocols.
@@ 4 Motivation: Synthetic Ridge Regression Experiment
We evaluate the Qwen 2.5 family (7B, 14B, 32B, 72B) (qwen2025qwen25technicalreport), across decoding temperatures $\{0.0,0.2,0.4,0.6,0.8\}$. Prompt templates are provided in Appendix A.3. To ensure valid outputs, generations are limited to 5 tokens with invalid predictions resampled.
@@ 5.1 Methodology
In practice, the Meta-Informed strategy assumes access to previous tasks along with high-performing configurations. For this study, we obtained such configurations by running extensive hyperparameter search with HEBO (Cowen-Rivers2022-HEBO) on a set of tabular regression and classification tasks. To maximize performance, ensembles (or blends) were built from the resulting models. We refer to the models with the highest contributions to these ensembles as Context Blends, and use them as the source of support examples passed to the prompt.
@@
The LLM is instructed to output a JSON object describing an ensemble of 10 models. For each supported family (CatBoost (Prokhorenkova2018), LightGBM, XGBoost (Chen2016), and scikit-learn MLP (Pedregosa2011)), we provide the model name, a list of valid hyperparameters, and a discrete grid of admissible values. This grid is included directly in the prompt, ensuring that the model generates configurations from a well-defined search space rather than free-form values. An excerpt of the schema is shown below (see Appendix F for full hyperparamater grids):
@@
We use the DeepSeek-R1 reasoning model (deepseekai2025deepseekr1incentivizingreasoningcapability), which naturally produces explanations of its choices. The LLM configuration is described in Appendix D. Invalid generations are rare, but we apply lightweight post-processing when they occur. If the LLM outputs a numeric value that falls outside the predefined hyperparameter grid, we project it to the nearest valid grid point. For non-numeric fields (e.g., categorical options) that cannot be matched, we discard the configuration and resample a fresh output. Likewise, if the JSON structure itself is malformed, the entire configuration is rejected and regenerated. Each run uses a different set of support examples, ensuring robustness to contextual variation.
@@
Each LLM call outputs 10 configurations, which we treat as candidate base models. We train these with cross-validation bagging and then combine their predictions using feedforward greedy blending (caruana2004ensemble). This procedure is applied consistently to LLM-based and baseline methods, providing a fair comparison and reflecting common ML ensembling practice.
@@ 5.3 Baselines
We compare LLM-based recommendations against four baselines representing different strategies for the CASH problem (full details in Appendix G): Context-Random (uniformly samples model–hyperparameter configurations from the same reference pool as the one passed to the LLM), Random-Hyperopt (at each step, uniformly samples a model family and then applies a hyperparameter optimizer within that family), LGBM-Hyperopt (optimizer restricted to LightGBM, capturing the strength of a single tuned family), and MaxUCB-Hyperopt (treats each family as a bandit arm, selecting the one with the highest upper-confidence bound before a single optimization step (balef2025cashbanditsmaxkarmed)). Context Blends consist of ensembles obtained from an extensive hyperparameter search. They provide upper-bound baselines: they achieve high performance through extensive search, and thus set the performance we seek to approach under a much more limited budget. All -Hyperopt baselines use HEBO (Cowen-Rivers2022-HEBO), chosen for its strong and consistent performance across diverse tasks (kegl2023systematicstudycomparinghyperparameter) [^1] HEBO begins with random search, using $1+\text{(dimension of the hyperparameter space)}$ evaluations, before switching to Bayesian optimization..
@@ A.1 Closed-Form Test Error
    Deterministic equivalents / resolvent fixed point. Using standard random-matrix techniques (resolvent identities and deterministic equivalents for sample covariance resolvents) [couillet2011random, Chapter 6], one shows that the random matrix inverse that appears in the ridge formula concentrates around the deterministic matrix $\bar{\mathbf{Q}}(\lambda)$ defined in equation 4 and that the scalar traces $(1/n)\operatorname{tr}(\mathbf{C}_{k}\bar{\mathbf{Q}})$ converge to the solution $\delta_{k}$ of equation 5. This gives the first-order deterministic equivalents used to compute $m_{k}$.
@@
    Gaussian fluctuation / CLT. After centering by the deterministic mean, the score is a linear or quadratic form of Gaussian vectors; a multivariate CLT (together with second-order deterministic equivalents captured by $\mathbf{K}_{k}$ and the $\mathbf{d}^{(j)}$ corrections) yields asymptotic Gaussianity tiomoko2020large with variance given by the deterministic formula $v_{k}$.
@@ Appendix F Base Model Details
We use four base models in our experiments: XGBoost [Chen2016], CatBoost [Prokhorenkova2018], LGBM, and SKMLP [Pedregosa2011]. The corresponding hyperparameter grids for each model are provided in Figure 6.
@@ G.4 MaxUCB-Hyperopt
For the MaxUCB-Hyperopt baseline, we implement the bandit-based CASH formulation proposed by balef2025cashbanditsmaxkarmed. In this setting, each candidate model family is treated as an arm in a multi-armed bandit, and hyperparameter optimization is carried out within the selected arm using HEBO. The Max-UCB algorithm balances exploration of new model families with exploitation of those that have already demonstrated promising performance.
@@ Appendix H Robustness to Prompt Shuffling
Large language models can sometimes exhibit position or recency biases [wang-etal-2023-primacy, wang2025eliminating], raising the question of whether the Meta-Informed strategy is sensitive to the way information is ordered inside the prompt. To test this, we generate two independent shuffled versions of the Meta-Informed prompt for each dataset–seed pair. In each shuffle, we randomly permute (i) the order of support datasets, (ii) the order of model families listed in the schema, and (iii) the order of hyperparameters within each family. The underlying content is unchanged, only the presentation order differs. The experimental setup is otherwise identical on the 22 Kaggle datasets the same contexts, ensembling pipeline, and $p_{\text{rank}}$ as the evaluation metric.